# The Role of Human Creativity in the Presence of AI Creativity Tools at Work: A Case Study on AI-Driven Content Transformation in Journalism

### v1

@@ 4.1.1. Reasons for AI Tool Usage
1. Selecting the narrative framing and premise. They all used the reframing features of ReelFramer to find the characters, premise, and key news facts in the web article that they could use to transform into a role-play script. They described this as a cognitively demanding task to “see the story” from a different narrative frame. ReelFramer’s suggestions were typically accurate, creating a reliable premise for a script.
@@
2. Drafting multiple scripts. They used the scriptwriting writing features of ReelFramer to initiate a draft of the script. However, the users consistently reported editing “about half” of the AI-generated script. They found they had to improve many of the writing mechanisms, such as taking out awkward phrases, increasing clarity, or removing lines with too little content. Overall, it was a helpful starting point.
On the other hand, none of the students used the storyboarding features, as they did not find this part of the task cognitively demanding. They said they could “see” how the story would be filmed and apparently did not use the prop, costume, or gesture suggestions—often because they were constrained by their personal set of potential props, costumes, and acting ability.
4.1.2. Reasons and Risks of Not Using the Tool
Creators chose not to use the AI tool for 6 of the 26 reels made during the study. There were several legitimate reasons not to use the tool. However, choosing not to use the tool sometimes resulted in problems.
1. Strong understanding of the initial news story. Two times, a participant who had used ReelFramer previously chose not to use the tool when making a news reel for a story they had reported. Because they conducted the initial research and interviews and framed the original written story, it was easier to understand the key points in the story and imagine it from multiple perspectives, making it cognitively easier to reframe around a role-play narrative.
2. Task deemed easy. P5 made his first news reel with ReelFramer but chose to make the other three without it because he understood the framework well enough to do it on his own. However, this decision had trade-offs. When the editor reviewed the reels made without ReelFramer, she appreciated several creative additions that the tool would not have suggested (Section 4.2). However, the scripts were insufficient for publication—as they all lacked sufficient information about the story. The task of creating news reels is often more difficult than it looks; balancing between information and entertainment is hard, and the tool seems to help balance those competing design requirements.
3. Independent inspiration. P2 typically used ReelFramer, but for his fourth video, he had an inspiration for a role-play narrative and developed the framing and script on his own. However, he deviated from the role-play framing and wrote the script as a monologue, which the editor and other creators found confusing. Once he iterated to return it to role-play form, it was published. This incident also shows the risks of not using the scaffolding provided by the tool, although this creator was able to recover.
4. AI narratives inappropriate for a serious topic. P1 typically used ReelFramer, but for her second video, ReelFramer was not producing a narrative that was serious enough for her topic, which dealt with domestic abuse. ReelFramer’s “quirky” and “fun” narratives clashed with the gravity of the topic. This clash led her to create the news reel on her own. However, she had written the initial news story, so she had a strong understanding of it, and she took inspiration from the tool and followed a role-play format that delivered three news facts. The editor approved it for publication, and this news reel received by far the most views. There could be many reasons for this success—topic, timing, luck, or unknown quirks of the platform’s sharing algorithm. This is discussed in Section 5. Despite her success in making a news reel without the tool, she chose to return to using ReelFramer for all her future videos due to the cognitive support it provided.
Generally, the group was successful in their task, producing 16 news reels garnering over 500,000 views. The tool was treated as a creative springboard, assisting with cognitively demanding tasks of reframing a narrative, and writing a script that balanced entertainment and information. However, creators did not need a creative springboard when they had independent inspiration or when they had a strong understanding of the original web article. Additionally, when the tool failed to provide an appropriate framing or script, they had to fall back on their own creativity.
Figure 3. Creative additions made by humans that AI does and does not support.
4.2. Role of Human Creativity: Human Creativity Where AI Left off
To understand when and how creators chose to use their own creativity rather than relying on AI, we analyzed all their creative additions made to news reels created during the study that ReelFramer did not (or could not) make. Every week, when the editor reviewed news reels made by the creators (both published and unpublished), she noted all the instances when the creator did something creative outside of what ReelFramer suggested. After the study, we analyzed the types of tasks creators choose to innovate on, and found three clear groups of times when creators made creative additions.
4.2.1. Human Creativity to Improve AI Outputs
In the study, the editor noted a few tasks that were supported by AI, but for which the creators decided to improve with their own creativity by using their own framing or scriptwriting choices.
For framing, there were two times when a creator rejected AI character suggestions, and came up with characters that AI would not have: using an inanimate object (a gas pump, to talk about oil prices) or introducing characters not in the story (an Uber Driver, and a passenger to talk about a new bridge). For scriptwriting, one creator merged facts from two different web articles about the same topic (a local election) to tell a better story. ReelFramer only supported creating scripts from a single web article, so the creator had to add details to the script herself.
Creators frequently had to edit AI’s attempts at humor in scripts. AI often used puns, which creators did not like. In the middle of scripts, creators typically removed the attempted humor, but at the end of scripts, creators needed to come up with a new ending like saying something sarcastic. One creator added a curse word for emphasis said by a construction worker character. AI would never have suggested a curse word, and although the editor made him bleep it out before publishing, the audience still understood the effect, and it was a bold and creative choice. Although editing humor was common, the other creative framing and scripting edits were rare—occurring in only 4 of the 26 news reels.
4.2.2. Human Creativity for Major AI Failures
Major AI failures were rare—occurring in 3 of the 26 videos, but required significant creativity on the part of creators. In all three cases, AI made socially inappropriate suggestions because it lacked context of the story or the creators.
The first story was about new doula services offered for expecting mothers. The original web story featured quotes and reactions from women of color, since they are disproportionately at risk during pregnancies. AI suggested roleplaying characters from the story who were pregnant Black women. However, the creators did not feel like they could play this in a respectful way. But since the service is generally available for all people, they changed the characters to be people indirectly affected by the service (concerned family) and were able to report on the effects for women of all races. By creatively changing the characters, the creators were able to reframe the story in a socially appropriate way.
The second story was about domestic violence, which was a subject that was too serious for the lighthearted tone in the AI-suggested script. The creator chose to write the script herself, although she was able to creatively adapt the role-play framing by making the Domestic Women’s Shelter the expert explaining the dangers of emergency alert testing to the “naive” FCC—emergency alerts can expose hidden phones that domestic abuse victims often have.
@@ 4.2.2. Human Creativity for Major AI Failures
The third, and most difficult case was when the team decided it was inappropriate to make a news reel out of a story. The article was titled “Women’s Club Basketball Discusses What Black History Month Means to Them”. None of the journalists identified as both Black and female, and thus role-play style narrative felt socially inappropriate. Additionally, they did not want to reframe the narrative around a different set of characters because it would displace the voices of the very people it was trying to center. Ultimately, they decided that not making the reel was the right decision.
Although major AI failures were relatively rare, they required human creativity to detect and address them.
4.2.3. Human Creativity Where AI Leaves off
In the study, the editor noted many tasks that were not supported by AI that creators decided to add their creativity to. This happened in 12 of 26 news reels (Figure 3). ReelFramer was designed to address the full creative process from framing to storyboarding. However, the creative process has no bounds, and the creators found many tasks to do creatively that the tool did not address.
There were four instances of creators adding personality to the characters, which was largely shown through acting ability. For example, making a teenager have an aggressively dismissive attitude about a new college program, to bring drama and realism to the character (as well as making it sound less like an advertisement).
There were three instances of using props in highly effective ways. First, using a real dog or baby instead of a doll, and second, in an episode about a soup kitchen, had a chef-like character actually making a dish and using a different cooking utensil in every scene to show progress. None of these were suggested by AI, but were things creators realized they had on hand.
The market research done in week 10 (Section 5) revealed that the videos would be more interesting if there were more movement, especially at the beginning. Thus, in the last week, creators added two more actions to videos – pouring maple syrup, jogging, and unwrapping food.
There were four other notable creative additions, like using animations of a truck for a story on monster trucks, adding sound effects, and adding colored lights and captions to reflect elements in a story.
Although the AI tool attempted to cover the entire creative process, creators found new dimensions where the tool left off, and they were able to add many of their own creative touches.
4.3. AI Risks: Need for Training Humans to Critique due to AI Overreliance
AI is widely known to produce errors—as are humans. The newsroom in this study had an editor who was in charge of quality control. This is typical of all newsrooms. Similar to typical editorial newsroom dynamics, creators submit drafts, receive critique and edits from the editor and have to iterate to produce a new draft. Iteration continues until the editor decides to publish the piece, or decide not to run it.
To understand the nature of the edits, we tracked two things: the magnitude of the error (major, minor or none), and the attribution of the error (whether the error was introduced by the AI or the human journalist). For error magnitude, major edits are those that required reshooting the entire news reel; minor edits required reshooting only part of the video (e.g., a localized scene or dialog edits); no edits meant it was ready for publication. Table 3 shows the magnitude of edits for each news reel’s first iteration.
We also tracked whether the errors were introduced by AI or by the creators, as judged by the editor from discussions with the creator. If the script contained incorrect information, and the script was drafted by AI and edited by a creator, the editor attributed this error to being introduced by AI. Although it should have been caught by a human, it might not have needed to be caught if it had not been introduced by AI. However, if the human writes a script with incorrect information, this error is attributed to the human.
Surprisingly, a vast majority of the errors were introduced by the AI tool. 11 of 15 major errors, and 15 of 19 minor errors were introduced by AI (Tables 1 and 2). These errors were accepted by the creators, and they filmed news reels with these errors in them. Although it is expected that AI will make errors, and that many people using AI on a deadline might simply accept AI content at face value, skimming the content and not checking it carefully. However, this case is different—creators actually had to speak the lines AI wrote, so they were fully cognizant of the dialog they were including. This was true for both student journalists and professional journalists (Section 6). It was surprising to the editor because being critical is an essential aspect of journalism training.
Major Issues (15) AI Attributed (11) Non-AI Attributed (4) too positive/promotional (4) slow paced (2) too long (2) confusing format - monologue (1) too few facts (2) biased narrative (1) unclear narrative (1) too wordy (1) abrupt intro (1)
@@ 4.3. AI Risks: Need for Training Humans to Critique due to AI Overreliance
Table 1. Major Issues: AI vs Non-AI Attributed issues in news reels created in the newsroom study.
Minor Issues (19) AI Attributed (15) Non-AI Attributed(4) awkward AI phrase (4) labels too long (1) confusing character names (3) added curse word (1) ending lacks polish (2) add relevancy to the present (1) too positive/promotional (1) missing facts (1) strays from news (1) tone shift (1) unclear/missing fact (2) unnecessary dialog (1)
@@
Table 2. Minor Issues: AI vs Non-AI Attributed issues in news reels created in the newsroom study.
@@
The most common major issue was clearly attributed to the AI tool—sounding too positive. The AI is trying to balance entertainment and information in a “quirky” way, and the stories should not sound sad and serious like traditional evening news anchors. However, the AI leaned too heavily toward a positive tone and this bias clashed with the content in different ways. For stories that were serious or sad (stories about homelessness, domestic violence, or climate change), the “quirky” tone was inappropriate for the topic. For many lighter stories—like a small town winning an award for being charming—AI’s positive tone was also a problem. The positive tone made the video feel like an advertisement rather than news. In all cases, the team had to find creative ways to fix the narratives to be more appropriate.
@@
The most common minor issue clearly attributed to the AI tool (and accepted by creators) was including bizarre and awkward AI phrases. These stick out badly and are fairly localized changes to make, so the effort to change them should have been minimal. Phrases like “a little bird told me’’ are awkward AI phrases that were inappropriate for a news reel because all news should make some effort to attribute information to sources. Other awkward phrases include “a comedy of wonders” and distracting attempts at wordplay such as “a two-tiered system? is that kind getting rewarded with two kinds of deserts?” When asked why they included these phrases, the typical response was “I guess it seemed OK.”
@@
This acceptance indicates an implicit trust in the AI, even when AI does a task in the creator’s realm of expertise. A crucial role of human creativity is to be critical of AI outputs. Even journalists who are trained to be critical may require additional training to learn to be critical in this new way.
4.4. Human Creative Role Over Time
We also tracked the magnitude and attribution of errors in the first iteration of all 26 news reels to track errors over time. As expected, every creator’s first video required major edits, but later videos needed more minor edits, or even no edits.
Participant Article 1 Article 2 Article 3 Article 4 Article 5 Article 6 P1 Major Minor Minor Minor No edits No edits AI errors 2 1 1 1 Non-AI errors 1 1 P2 Major Minor Major Major Major Minor AI errors 3 1 3 1 Non-AI errors 1 1 P3 Major Minor Minor Major / / AI errors 3 2 2 2 / / Non-AI errors / / P4 Major Minor Major Minor Minor No edits AI errors 1 1 2 1 1 Non-AI errors P5 Major Major Major Minor / / AI errors 2 / / Non-AI errors 1 1 1 1 / /
Table 3. Classification of the amount of editing needed for news reels made for each article and each participant. Edits were either major (requiring a full reshoot), minor (localized scene or line edits) or no edits. Errors are broken into those introduced by AI (and accepted by creators without realizing), or errors not introduced by AI. Cells filled with a slash indicate creators did not make news reels (some participants made more than others).
4.4.1. Learning the Limits of the Tool, and Reasserting Professional Standards

### v2

@@ 3.2.1. Reasons for Usage: Framing and Drafting Scripts
1. Narrative framing. They all used the narrative framing features of ReelFramer to find the characters, premise, and key news facts in the web article that they could use to transform into a role-play script. They described this as a cognitively demanding task to “see the story” from a different narrative frame. ReelFramers’ suggestions were typically accurate, creating a reliable premise for a script.
2. Drafting scripts from premises. They used the scriptwriting features of ReelFramer to initiate a draft of the script. Although generating the premise was helpful, there was another cognitive challenge of expanding that to a script. Translating raw ideas into fluent text is a complex writing task that requires perspective-taking, thematic grouping and sequencing of information, and managing cohesion and flow  citeflower1981cognitive.
@@
However, the users consistently reported editing “about half” of the AI-generated script. They found they had to improve many of the writing mechanisms, such as taking out awkward phrases, increasing clarity, or removing lines with too little content. Overall, it was a helpful starting point.
Notably, those who used the tool only used it for the narrative framing and scripting steps. None of the students used the storyboarding features, as they did not find this part of the task cognitively demanding; they said they could “see” how the story would be filmed.
3.2.2. Reasons (and Risks) of Disuse
Creators chose not to use the AI tool for 6 of the 26 reels. In some cases, this was reasonable; in others, it led to problems.
1. Strong understanding of the story (2). Two creators skipped ReelFramer when adapting stories they had originally reported. Having done the initial reporting made it easier to reframe the story into a role-play narrative without the tool.
2. Task deemed easy (2). P5 used ReelFramer once, then stopped, believing he understood the framework. While his reels had creative additions and were entertaining, they were not publishable as they all lacked key news information. This suggests ReelFramer could help as a scaffolding tool, even for those who do not want or need it.
3. Independent inspiration (1). P2 had an independent inspiration for his own framing (a monologue rather than a role-play), which he executed by himself. The editor liked the inspiration, but found the monologue format confusing. He iterated to bring back the role-play style framing and was ultimately successful in publishing the video. Independent inspiration could become more scaffolding.
4. Tone-deaf AI (1). P1 skipped the tool when its “fun” and “quirky” tone did not match the seriousness of the topic—domestic violence. She took inspiration from the format she had learned from the AI tool and wrote her own script, and the video was successfully published. For her future reels, she returned to the tool due to its cognitive support.
3.2.3. Editors’ Reflections Over Time
In general, the tool made two cognitively challenging tasks easier—it was not always necessary, but it was helpful. The editor observed that in the beginning, the tool served as a teaching instrument to scaffold the process during the learning phase. This was particularly helpful for reframing—whereas journalists are taught to write stories in a style that is typically serious, and written from a more objective point of view, the reframing steps that help journalists “see” a story in this new way, also provided the scaffolding to help them find these transformations for themselves.
Additionally, when a user chose not to use the tool, their reels had fundamental problems such as not balancing information with entertainment—the core challenge the tool supports. For some users, positioning the tool only as a scaffolding tool for learning could help them understand the complexities of the task before setting out on their own.
There were incidents when the tool failed, and it was clear that the creators could make new reels without it. However, 4 of the 5 users still chose to use the tool. They treated it as a “creative springboard”—rather than a replacement for their creativity, particularly when it came to drafting scripts based on the premise. Finding the words to use in a script is hard, and it is easier to edit a script than generate one from scratch.
3.3. Findings: Role of Human Creativity
Figure 3. Creative additions made by users.
To understand when and how creators chose to use their own creativity rather than relying on AI, we analyzed all their creative additions made to news reels created during the study that ReelFramer did not (or could not) make. This analysis is based on a list of creative additions noted by the editor every week during feedback sessions. This includes both published and unpublished reels. Some changes, like correcting an awkward AI phase, or adding captions, or correcting facts, were not deemed creative, as they were fairly typical copy editing changes. However, changes like introducing new types of characters, or adding new elements (actions, personality, animations, etc) were deemed creative because they significantly deviated from the outputs that AI could have produced.
Across the 26 videos made by the students, the editors noted 22 instances where the creator made substantial creative additions (see Figure 3). The editor also denoted the four parts of the creative process where they were made.
1. Framing or Premise (7) Overall, 7 creative additions were made in the framing or premise of the news reel. Whereas framing and premise are supported by the AI tool, users choose to make improvements to this area based on either inspiration or need. This includes finding different characters to base the role play around, introducing new structures for the narrative, and deciding not to make a news reel at all.
Whereas the tool limited itself to using characters directly taken from the news story, the users knew that this was not strictly necessary for TikTok videos. They could use inanimate objects or tangential characters if they were more fun or easier to depict in the video. The tool was also artificially limited in its structures, but users needed to change that—this included merging facts from two news stories to improve the narrative (rather than working with only one story as the tool assumes). The tool also assumes the tone should be quirky and fun, but for a story about a serious subject of domestic violence, the creator realized the “fun” tone was inappropriate, and decided to make a role-play but with a serious tone. This was a large departure from the tool—and the typical newsreel format—but it was necessary, and the reel got the most views by far.
@@ 3.3. Findings: Role of Human Creativity
Users also needed to exercise creative and professional judgment to determine when a news reel format would be inappropriate. The article was titled “Women’s Club Basketball Discusses What Black History Month Means to Them”. None of the journalists identified as both Black and female, and thus role-play narrative felt socially inappropriate. Additionally, they did not want to reframe the narrative around a different set of characters because it would displace the voices of the people it was trying to center. Ultimately, they decided that not making the reel was the right decision.
2. Script (2) Overall, most users appreciated AI most for its ability to produce scripts. All AI-generated scripts needed editing, but sometimes those edits had to be creative. AI would often add terrible jokes to scripts and users had to find a way to improve the humor—sometimes they deliver the AI joke sarcastically, to make it funny (which is not how it was intended). Sometimes, lines were dull and needed more energy. One user added a curse word, which the editor insisted he bleep, but it still brought more energy and was funny. In both cases, users brought their own sense of humor to the script.
3. Reel Production (10) A major area where users added creativity was in the production of the reel—acting, props, lighting, sound effects, and post-production like actions and animation. This is an area that the tool did not support well, and that users found there were many ways to enhance the quality of reels—as well as bring in their own knowledge and creativity. Adding a strong personality to the characters was a creative decision that made a big difference. AI scripts had a negative tendency to sound too positive or even promotional. But creators could take an enthusiastic character and make them sarcastic, or have them be rudely cut off by another character to undermine the overzealous positivity. Giving characters amplified emotions (such as being very surprised, or very miserable when sick) also gave videos more life.
Using props well also made a big difference, and creators used what they had on hand—including a real dog for a story about animal shelters, or cooking utensils for a story about food insecurity. AI could not have known what users would have, so its suggestions for props were weak, but users reacted to their environment to add drama through prop usage. Additionally, CapCut software makes post-production effects like animations, sound effects, lighting filters, and caption colors very easy. Users often found creative ways to exchange the message through this type of post-production.
4. Iterating from Audience Feedback (3) In the last week, creators added three more actions to videos—pouring maple syrup, jogging, and unwrapping food. Each of these was highly relevant to the story and made the videos more fun and eye-catching.
Overall, users made significant creative additions to their reels. This included adding to processes the tool supported, such as framing and scripting, as well as places where AI left off—such as acting, visual effects, and props, or iterating based on user feedback. When making these additions, users often injected their own knowledge and taste, like their own sense of humor, their acting ability, their knowledge of post-production techniques, and the props they had on hand. Moreover, this level of creativity indicates that users were not passively accepting AI’s creativity. They actively inserted themselves in the process based on need (when AI failed) or inspiration (when they had independent ideas—often that were better than AI’s suggestions).
3.3.1. Editors’ Reflections Over Time
To the editor, it was clear that the creators were heavily invested in producing high-quality videos. Whereas AI could produce many outputs, the creators took ownership over the entire process. They critiqued outputs, considered the audience, adjusted based on what was appropriate and available, and generally tried to maximize the creative potential of each story, within the time constraints of their deadline. In stark contrast, AI was mostly following a template and executing a task as directed. The creators’ sense of ownership and holistic responsibility for the quality of the product drove their continued creativity.
3.4. Overcoming AI Errors
Major Issues (15)
AI Attributed (11) Human Attributed (4)
too positive/promotional (4)
too long (2)
too few facts (2)
unclear narrative (1)
too wordy (1)
abrupt intro (1)
slow paced (2)
confusing format – monologue (1)
biased narrative (1)
@@ 3.4. Overcoming AI Errors
Table 1. Major Issues: AI vs Human Attributed in the student newsroom study.
Minor Issues (19)
AI Attributed (15) Human Attributed (4)
awkward AI phrase (4)
confusing character names (3)
ending lacks polish (2)
too positive/promotional (1)
strays from news (1)
tone shift (1)
unclear/missing fact (2)
unnecessary dialogue (1)
labels too long (1)
added curse word (1)
add relevancy to the present (1)
missing facts (1)
@@
Table 2. Minor Issues: AI vs Human Attributed in the student newsroom study.
AI is widely known to produce errors—as do humans. The newsroom in this study had an editor who was in charge of quality control. This is typical of all newsrooms. Similar to typical editorial newsroom dynamics, creators submit drafts, receive critiques and edits from the editor, and have to iterate to produce a new draft. Iteration continues until the editor decides to publish the piece or decides not to run it.
To understand the nature of the edits, we tracked two things: the magnitude of the error (major or minor) and the attribution of the error (whether the error was introduced by the AI or the human journalist).
For error magnitude, major edits are those that required reshooting the entire news reel; minor edits required reshooting only part of the video (e.g., a localized scene or dialog edits). We also tracked whether the errors were introduced by AI or by the creators, as judged by the editor from discussions with the creator. If the script contained incorrect information, and the script was drafted by AI and edited by a creator, the editor attributed this error to being introduced by AI. Although it should have been caught by a human, it might not have needed to be caught if it had not been introduced by AI. However, if the human writes a script with incorrect information, this error is attributed to the human.
Surprisingly, a vast majority of the errors were introduced by the AI tool. 11 of 15 major errors, and 15 of 19 minor errors were introduced by AI (Tables 1 and 2). These errors were accepted by the creators, and they filmed news reels with these errors in them. Although it is expected that AI will make errors, and that many people using AI on a deadline might simply accept AI content at face value and not check it carefully. However, this case is different—creators actually had to speak the lines AI wrote, so they were fully cognizant of the dialogue they were including.
@@
The most common major issue was clearly attributed to the AI tool—sounding too positive. The AI is trying to balance entertainment and information in a “quirky” way, and the stories should not sound sad and serious like traditional evening news anchors. However, AI leaned too heavily toward a positive tone, and this bias clashed with the content in different ways. For stories that were serious or sad (stories about homelessness, domestic violence, or climate change), the “quirky” tone was inappropriate for the topic. For lighter stories—like a small town winning an award for being charming, AI’s positive tone was also a problem. The positive tone made the video feel like an advertisement rather than news. In all cases, the team had to find creative ways to fix the narratives to be more appropriate.
@@
The most common minor issue clearly attributed to the AI tool (and accepted by creators) was including bizarre and awkward AI phrases. These stick out badly and are fairly localized changes to make, so the effort to change them should have been minimal. Phrases like “a little bird told me” are awkward AI phrases that were inappropriate for a news reel because all news should make some effort to attribute information to sources. Other awkward phrases include “a comedy of wonders” and distracting attempts at wordplay such as “a two-tiered system? Is that kind getting rewarded with two kinds of desserts?” When asked why they included these phrases, the typical response was “I guess it seemed OK.”
@@
This acceptance indicates an implicit trust in AI, even when the AI does a task in the creator’s realm of expertise. A crucial role of human creativity is to be critical of AI outputs. Even journalists who are trained to be critical may require additional training to learn to be critical in this new way.
3.4.1. Editors’ Reflection Over Time
The editor noticed a general trend that, initially, users relied heavily on the AI’s outputs—even accepting its blatant errors without question. However, through critical reflection and iteration, creators were able to reassert themselves. AI was initially seen as an authority on news reels, due to the creators’ lack of experience and AI’s mystique. But ultimately, users were better able to balance AI’s authority and their own.